i need help for page 96. the info you requested is on page 95.

Cveni hass Tablet3-0ea5 iet Sample 0.a4764 Somal Tanes 23. 9 Pessue125-0 teir -o.95894 -9539 A QUESTIONS 10. Using your answer to problem 1, the theoretical per cent of HCO, in an alka-seltzer tablet, answer the following. A student weighed an alka-seltzer tablet at 3.0625g, He then broke a sample weighing 0,2478 g from that tablet to run the exneriment exactly as it was run in this lab. The room temperature was 23.0°C. Atmospheric pressure was measured as 725.0 torr. The student used exactly 5.00 mL of HCI to run the reaction. a. Use dimensional analysis and the theoretical per cent bicarbonate to calculate the number of g HCO in the sample. 3. blea5 x 43.955 HCo = 1. 3459e Tablet 100 Table 1.34(oC b. Use dimensional analysis to calculate the maximum number mole CO2 which would be generated from the sample. 5HCC3 hol HCOs nol COa c. Use Dalton's law and the vapor pressure of water at this temperature to determine the pressure of the CO; in the apparatus in atmospheres. Pr Pcost PH d. Use the ideal gas law to calculate the volume in L that the CO2 generated from the sample is expected to occupy. Convert your answer to mL. The student became concerned because the volume observed in the gas buret was 4.25 mL lower than what was calculated. Should the student have been concerned? Why or why not? Ven RT V-nRT Pcoa P 96 Page
B. Room temperature 3.3aa C. Mass of tablet D. Mass of NaHCO, in the tablet (from label) .9/l0 no leaks DATA TABLE Trial 3 Mass of sample (g) O-a1l5 Sa .a54s e 31-20mL 1. 2.Vol (final mL) 3160かL 27.00mL to OmL 3.Vol (initial mL) OmL 3700aL +0C mL 4.Vol (gas mL): (Vfnal- Vinitial) 3)00 mし 31. T0ML a7L 5.Constant: 270U mLt 4: 8ULt4. 30Lt400 35.00m 35-70ML hL 4.00 mL CO2 absorbed per 5.00 mL HC ant t 31 6.Vol.co2 oal (Vgas+ constant) 07-4 tos 58 CALCULATIONS 1. From the mass of NaHCO) in the tablet (D). calculate the mass in grams of HCO in the tablet. Use dimensional analysis. 1.4/p9 Na HCOS ? HCO low: 1.91 9 NaHCO Inu IANH ARHEO mat He 400 HCO a4sHcO 4598 2の代NSH8
Answer #1

#10:(a): Mass of sample = 0.2478 g

Theoretical % of bicarbonate = 43.95%

=> g HCO3- in the sample = 0.2478 g * (43.95 / 100) = 0.10891 g

(b): moles of HCO3- in the sample = mass/molar mass = 0.10891 g / 61.02 g/mol = 0.0017848 mol

1 mol HCO3- releases 1 mol CO2. Hence mole ratio is 1:1

Hence maximum number of moles of CO2 generated from the sample = 0.0017848 mol

(c): Vapor pressure of water at 23.0 oC = 21.1 torr

Pt = 725.0 torr = P(CO2) + P(H2O)

=> 725.0 torr = P(CO2) + 21.1 torr

=> P(CO2) = 725.0 - 21.1

=> P(CO2) = 703.9 torr

P(CO2) in atmosphere = 703.9 torr * (1 atm / 760 torr) = 0.9262 atm

(d): Applying ideal gas equation:

V = nRT/P

=> V = 0.0017848 mol * 0.08206 L.atm.mol^-1K^-1 * (273.15+23.0) / 0.9262 atm

=> V = 0.04683 L

Volume in mL = 0.04683 L * (1000 mL / 1L) = 46.83 mL